### v1

@@ 3.3 Integration
We are concerned with the assembly of $\boldsymbol{A}$, $G_{i}$, and $\frac{\partial G_{i}}{\partial\boldsymbol{c}_{j}}$
  1. 1.
    $\boldsymbol{A}=\left\langle\frac{\partial\Phi^{N}}{\partial\omega}\cdot s,\Phi^{N}\right\rangle_{\Omega}$: Each component of this matrix is computed using the calculation
[…]
    where $s(\omega)=\left(s_{1}(\omega),\dots,s_{d}(\omega)\right)^{\top}$. Depending on the expansion of each $s_{k}(\omega)$, for example if they are monomials, this expression can be computed exactly.
  2. 2.
    $G_{i}(\boldsymbol{c})=-\left\langle f_{i}(\pi^{N}(\omega),\ell(\omega)),\Phi^{N}\right\rangle_{\Omega}$: Each component of the vector function $G$ is represented as
[…]
    The evaluation of these integrals for $i=1,\dots,n$ can be approximated using a suitable quadrature rule, for example the Gauss-Legendre quadrature in $d$ dimensions over the domain $\Omega$, up to a desired accuracy.
  3. 3.
    $\frac{\partial G_{i}}{\partial\boldsymbol{c}_{j}}=-\left\langle\frac{\partial f_{i}(\pi^{N}(\omega),\ell(\omega))}{\partial\boldsymbol{c}_{j}}\Phi^{N},\Phi^{N}\right\rangle_{\Omega}$: We proceed similarly as for $G_{i}(\boldsymbol{c})$. We have the expression
    (39) $\frac{\partial G_{i}}{\partial\boldsymbol{c}_{j}}_{k,l}$ $=-\left\langle\frac{\partial f_{i}(\pi^{N}(\omega),\ell(\omega))}{\partial\boldsymbol{c}_{j}}\phi_{l}\phi_{k}\right\rangle_{\Omega}$
[…]
    This integral is approximated by a suitable quadrature rule.
@@ 3.4 Computational complexity and implementation
The bulk of the computation comes from integration over $d$ dimensions, as well as the inversion of the matrix $JF$. The computational complexity of integration over $d$ dimensions scales exponentially with $d$, meaning that if the nonlinearities in $f$ do not allow a convenient matrix or tensor representation of $G_{i}$ as discussed in Remark 3.2, then we encounter the curse of dimensionality in computing the integral when $d$ becomes large.
The inversion of the Jacobian matrix scales approximately cubically with the size of the matrix, $Nn$. This is therefore dependent on the size of the system, $n$, and is indirectly combinatorially dependent on the dimensionality $d$ of the signal generator and the maximum degree $M$. Part of the objective of our polynomial approximation is therefore to find an appropriate, relatively low, maximum degree $M$ where $\pi^{N}$ sufficiently approximates the solution to (16), and to see how the approximation error scales with the maximum degree.
3.5 Evaluating approximate solutions
The approximate solution $\pi^{N}(\omega)$ can be verified by analyzing the residual of the equation (16). In essence, we are measuring the residual error of $n$ different PDEs, and so we can define the residual error function as
(40) $R(\boldsymbol{c},\omega)=(R_{1}(\boldsymbol{c},\omega),\dots,R_{n}(\boldsymbol{c},\omega))^{\top},$
where for each dimension $i=1,\dots,n$ the residual at a specified $\omega$ is given by
(41) $R_{i}(\boldsymbol{c},\omega)=\frac{\partial\pi_{i}^{N}}{\partial\omega}s(\omega)-f(\pi^{N}(\omega),\ell(\omega)).$
In order to quantify how well a given set of coefficients $\boldsymbol{c}$ approximates the solution, we compute the $L_{2}$ norm of each $R_{i}$, averaged over a chosen subdomain $W\subseteq\Omega$, as
[…]
Finally, to generate a single metric measuring the success of the approximation, we consider the average over all $i=1,\dots n$ of $\lVert R_{i}\rVert_{(2,W)}$, weighted by the relative magnitude of the coefficients $\boldsymbol{c}_{i}$:
@@ 3.5 Evaluating approximate solutions
4 Numerical results
Different numerical tests were performed to assess the domain over which a good approximation could be found with a relatively low maximum degree in the Galerkin expansion, due to the increasing computational complexity of the problem with higher dimensional expansions. These tests were carried out over chosen rectangular domains $\Omega$ of varying sizes, for different nonlinear dynamics and signal generators, and the algorithm was employed with a tolerance of $10^{-7}$ in the $L^{1}$ norm of $F$. Two Python modules for solving the invariance PDE and for generating the reduced-order model, as well as several scripts to replicate the results shown in this section, can be found at https://github.com/carlosjdoebeli/Moment-Matching-ROM.
4.1 Summary of examples
We primarily looked at four different examples to test our method and its limitations. The first two examples serve as a proof-of-concept for the method and shows its ability to recover known analytical solutions from low-dimensional nonlinear systems. The first two tests, outlined in Section 4.2 and Section 4.3, are nonlinear, and demonstrate the algorithm’s ability to recover highly accurate solutions when the components of $\pi$ belong to the span of the basis.
The third and fourth examples, in Section 4.4 and Section 4.5, show the method’s ability to recover solutions for larger nonlinear systems with both a linear and nonlinear signal generator. Tests are done with a full-order model dimensionality of $n=2$, $n=100$, and $n=1000$, to demonstrate the scalability of the method to large nonlinear systems. Because the analytical solution is not known, we test the reduced-order model’s ability to recreate the signal produced by the full-order model in these cases, alongside evaluating the residual error.
4.2 Test 1: Low-dimensional nonlinear system
First, we consider the model presented in [20, Example 8.1.3] with $n=2,d=2$ given by
$s(\omega)$ $=a\begin{bmatrix}\omega_{2}\\ -\omega_{1}\end{bmatrix}$ $\ell(\omega)$ $=\omega_{1}$
$f(x,u)$ $=\begin{bmatrix}-x_{1}+u\\ -x_{2}+x_{1}u\end{bmatrix}$ $h(x)$ $=x_{1}.$
For this problem, the map $\pi$ admits the analytical solution
$\pi(\omega_{1},\omega_{2})=\left(\frac{1}{1+a^{2}}(\omega_{1}-a\omega_{2}),\ \frac{1}{1+5a^{2}+4a^{4}}((1+a^{2})\omega_{1}^{2}-3a\omega_{1}\omega_{2}+3a^{2}\omega_{2}^{2})\right)^{\top}.$
Note that the answer can be expressed in monomials in $\omega$ up to degree two, which lies in the span of our chosen basis.
Using our representations of $G_{i}$ in (38) and $\frac{\partial G_{i}}{\partial\boldsymbol{c}_{j}}$ in (39), we can compute the function $F$ and its Jacobian as
$F(\boldsymbol{c})=\begin{bmatrix}\boldsymbol{A}\boldsymbol{c}_{1}+\boldsymbol{M}\boldsymbol{c}_{1}-\boldsymbol{\gamma}\\ \boldsymbol{A}\boldsymbol{c}_{2}+\boldsymbol{M}\boldsymbol{c}_{2}-\boldsymbol{P}\boldsymbol{c}_{1}\end{bmatrix},\quad JF(\boldsymbol{c})=\begin{bmatrix}\boldsymbol{A}+\boldsymbol{M}&0\\ -\boldsymbol{P}&\boldsymbol{A}+\boldsymbol{M}\end{bmatrix}$
using the previously defined matrix $\boldsymbol{A}$ in (37), as well as defining
(44) $\boldsymbol{M}=\left\langle\Phi^{N},\Phi^{N}\right\rangle_{\Omega},$ $\boldsymbol{P}=\left\langle\Phi^{N}\omega_{1},\Phi^{N}\right\rangle_{\Omega},$ $\boldsymbol{\gamma}=\left\langle\ell(\omega),\Phi^{N}\right\rangle_{\Omega}.$
Using a constant $a=2$ and a domain of $\Omega=[-1,1]^{2}$, the Galerkin method attains nearly exact recovery of the analytical solution, as shown in Table 1.
Domain $M=2$ $M=4$ $M=6$
$\Omega=[-1,1]^{2}$ $1.1048\times 10^{-16}$ $1.0940\times 10^{-15}$ $5.7176\times 10^{-14}$
Table 1: Table of the average residual error $\lVert R(\boldsymbol{c},\omega)\rVert_{(2,W)}$ for the Galerkin approximation of the invariant mapping $\pi$ for Test 1 with a constant $a=2$. The error is shown for different values of $M$ over the domain $W=[-1,1]^{2}$.
4.3 Test 2: Cart pendulum
We also consider the position control of a nonlinear cart pendulum, based on a problem studied in [32, Section 3.2]. The signal generator is chosen by selecting target dynamics for the cart pendulum with $d=2$ and $n=4$. The dynamics of the interconnected system are then governed by the equations
@@ 4.3 Test 2: Cart pendulum
This problem is known to admit an analytical solution given as
@@
With respect to our algorithm, $F$ and $JF$ are represented as
$F(\boldsymbol{c})=\begin{bmatrix}\boldsymbol{A}\boldsymbol{c}_{1}-\boldsymbol{M}\boldsymbol{c}_{3}\\ \boldsymbol{A}\boldsymbol{c}_{2}-\boldsymbol{M}\boldsymbol{c}_{4}\\ \boldsymbol{A}\boldsymbol{c}_{3}-H(\boldsymbol{c}_{1})\\ \boldsymbol{A}\boldsymbol{c}_{4}-\boldsymbol{\beta}\end{bmatrix},\quad JF(\boldsymbol{c})=\begin{bmatrix}\boldsymbol{A}&0&-\boldsymbol{M}&0\\ 0&\boldsymbol{A}&0&-\boldsymbol{M}\\ -\frac{\partial H}{\partial\boldsymbol{c}_{1}}&0&\boldsymbol{A}&0\\ 0&0&0&\boldsymbol{A}\end{bmatrix}$
where $\boldsymbol{A}$, $\boldsymbol{M}$ and $\boldsymbol{\gamma}$ are defined as before in (37), and (44), and the function $H:\mathbb{R}^{N}\to\mathbb{R}^{N}$ is given by
@@
In this problem, due to the choice of $s(\omega)$, the matrix $A$ is singular, and therefore the Jacobian $JF$ is also singular. Therefore, we must use the pseudoinverse to solve for our optimal $\pi^{N}$. We must then be careful in selecting our starting guess for $\boldsymbol{c}^{0}$; since the analytical solution contains coefficients that are mostly zero, we choose a starting guess of constants that are uniformly zero.
Using the constants $a_{1}=2$, $a_{2}=3$, and $k=-\frac{2}{3}$, and a starting guess of $\boldsymbol{c}=0$, the Galerkin method is able to recover the analytical solution. Table 2 once again shows a negligible error in the solution, close to machine precision, and reducing the tolerance yields even smaller errors. Note that in both this and the previous example, the solution admitted a monomial representation and the algorithm was able to achieve exact recovery.
Domain $M=2$ $M=4$ $M=6$
$\Omega=[-1,1]^{2}$ $9.3256\times 10^{-10}$ $8.6338\times 10^{-10}$ $9.4227\times 10^{-10}$
Table 2: Table of the average error $\lVert R(\boldsymbol{c},\omega)\rVert_{(2,W)}$ for the Galerkin approximation of the invariant mapping $\pi$ for Test 2 with a constants $a_{1}=2,a_{2}=3,k=-\frac{2}{3}$. The error is shown for the domains $\Omega=[-1,1]^{2}$, as well as for different maximal degrees of $m=3$, $m=4$, and $m=6$.
4.4 Test 3: Nonlinear resistor-inductor ladder with linear oscillator as signal generator
We consider the nonlinear resistor-inductor (RL) ladder, based on a slightly modified version of the schematic in [34, Figure 7] and also appearing in a similar context for model order reduction in [24]. The signal generator has dimension $d=2$, and the system has a repeating structure that allows for considering any number of dimensions $n$. This experiment is particularly interesting because there is no analytical solution, and it will allow us to determine whether our method is able to find a good approximate solution. The signal generator is given by:
@@ 4.4 Test 3: Nonlinear resistor-inductor ladder with linear oscillator as signal generator
The dynamics above depend on a parameter $\kappa$. Using $\kappa=1$, the system no longer satisfies the necessary assumptions above a certain number of degrees $n$, and becomes unstable, so we use a value of $\kappa$ close to 1 but slightly bigger than 1. Since the nonlinearities in each $f_{i}$ are polynomials of degree 2 and 3 and only depend on the corresponding $x_{i}$, we can make the following simplifications in the construction of $F$ and $JF$.
We begin by defining the tensors $\boldsymbol{N}$ and $\boldsymbol{O}$:
[…]
We now also introduce the functions $\tilde{N}:\mathbb{R}^{N}\to\mathbb{R}^{N\times N}$ and $\tilde{O}:\mathbb{R}^{N}\to\mathbb{R}^{N\times N}$, intended to act upon each $\boldsymbol{c}_{i}$. They are defined by
(48) $\left[\tilde{N}(\boldsymbol{v})\right]_{i,j}=\sum_{k=1}^{N}v_{k}N_{i,j,k},$ $\left[\tilde{O}(\boldsymbol{v})\right]_{i,j}=\sum_{k=1}^{N}\sum_{l=1}^{N}v_{k}v_{l}O_{i,j,k,l}.$
Finally, define the functions $P,Q:\mathbb{R}^{N}\to\mathbb{R}^{N\times N}$ as
(49) $P(\boldsymbol{v})=\left(\boldsymbol{A}+2\kappa\boldsymbol{M}+\frac{1}{2}\tilde{N}(\boldsymbol{v})+\frac{1}{3}\tilde{O}(\boldsymbol{v})\right)\boldsymbol{v}$
$Q(\boldsymbol{v})=\boldsymbol{A}+2\kappa\boldsymbol{M}+\tilde{N}(\boldsymbol{v})+\tilde{O}(\boldsymbol{v}),$
where $\boldsymbol{A}$ and $\boldsymbol{M}$ are as given in (37) and (44).
Then the function $F$ is represented in each dimension as
@@
where $\boldsymbol{\gamma}$ is the same as in (44), and its Jacobian is given by
@@
The construction of $F$ through these tensors, which are computed exactly rather than by numerical integration, allows us to significantly reduce the computational complexity of the problem by taking advantage of the system’s dynamics. We compute the approximate solution $\pi^{N}$ first for a low-dimensional full-order model with $n=2$, and then with high-dimensional full-order models with $n=100$ and $n=1000$. In each of these cases, we test the solution over different domains $\Omega=[-1,1]^{2}$, $\Omega=[-2,2]^{2}$, and $\Omega=[-3,3]^{2}$, as well as for different maximum degrees $M=2$, $M=4$, and $M=6$. The error in each case is measured over the same subdomain $W=[-0.7,0.7]$.
The resulting average residual error $\lVert R(\boldsymbol{c},\omega)\rVert_{(2,W)}$ for $n=2$ is shown in Table 3. In general, the error decreases with an increased maximum degree of the approximation, as the class of solutions for lower maximum degrees is contained in the class of solutions for higher degrees. The residual error increases when the domain $\Omega$ is enlarged for the same degree of expansion. However, for the larger domains, the method still recovers a relatively good approximation of the mapping $\pi$, implying that the region $W^{0}$ about the origin where the invariant is defined contains the rectangle $[-3,3]$.
Domain $M=2$ $M=4$ $M=6$
$\Omega=[-1,1]^{2}$ $1.3626\times 10^{-3}$ $1.9130\times 10^{-5}$ $7.0587\times 10^{-7}$
$\Omega=[-2,2]^{2}$ $8.2881\times 10^{-3}$ $4.7759\times 10^{-4}$ $7.9316\times 10^{-5}$
$\Omega=[-3,3]^{2}$ $2.1234\times 10^{-2}$ $1.3698\times 10^{-3}$ $7.1681\times 10^{-4}$
Table 3: Table of the average residual error $\lVert R(\boldsymbol{c},\omega)\rVert_{(2,W)}$ over the subdomain $W=[-0.7,0.7]^{2}$ for different values of $\Omega$ and $M$ for the Galerkin approximation to the invariant mapping $\pi$ with $n=2$ for the RL ladder problem with a linear oscillator signal generator with constants $a=2$ and $\kappa=1.1$.
The corresponding results and residual error $\lVert R(\boldsymbol{c},\omega)\rVert_{(2,W)}$ for $n=100$ are shown in Table 4. Once again, we can see that the method recovers a similarly good approximation of the solution. The error is on the same order of magnitude as it is for $n=2$, and is even slightly smaller in most cases. All tests are able to recover the mapping quite well with a maximum polynomial degree of 6. We also show the results for $n=1000$ in Table 5. We see similar results for $n=100$ and $n=1000$ for this test. This example demonstrates that the algorithm is able to recover the solution to 1000 dimensions for the RL ladder circuit with a linear oscillator as a signal generator.
Domain $M=2$ $M=4$ $M=6$
$\Omega=[-1,1]^{2}$ $1.0617\times 10^{-3}$ $2.2500\times 10^{-5}$ $7.0723\times 10^{-7}$
$\Omega=[-2,2]^{2}$ $6.5405\times 10^{-3}$ $5.8132\times 10^{-4}$ $8.1913\times 10^{-5}$
$\Omega=[-3,3]^{2}$ $1.7266\times 10^{-2}$ $1.7815\times 10^{-3}$ $7.6927\times 10^{-4}$
Table 4: Table of the average residual error $\lVert R(\boldsymbol{c},\omega)\rVert_{(2,W)}$ over the subdomain $W=[-0.7,0.7]^{2}$ for different values of $\Omega$ and $M$ for the Galerkin approximation to the invariant mapping $\pi$ with $n=100$ for the RL ladder problem with a linear oscillator signal generator with a constants $a=2$ and $\kappa=1.1$.
Domain $M=2$ $M=4$ $M=6$
$\Omega=[-1,1]^{2}$ $1.0617\times 10^{-3}$ $2.2500\times 10^{-5}$ $7.0723\times 10^{-7}$
$\Omega=[-2,2]^{2}$ $6.5405\times 10^{-3}$ $5.8132\times 10^{-4}$ $8.1913\times 10^{-5}$
$\Omega=[-3,3]^{2}$ $1.7266\times 10^{-2}$ $1.7815\times 10^{-3}$ $7.6927\times 10^{-4}$
Table 5: Table of the average residual error $\lVert R(\boldsymbol{c},\omega)\rVert_{(2,W)}$ over the subdomain $W=[-0.7,0.7]^{2}$ for different values of $\Omega$ and $M$ for the Galerkin approximation to the invariant mapping $\pi$ with $n=1000$ for the RL ladder problem with a linear oscillator signal generator with a constants $a=2$ and $\kappa=1.1$.
To give a representative example of the computational cost of the algorithm, Table 6 shows the amount of time required for the algorithm to converge for $\Omega=[-1,1]^{2}$ and $M=6$ for the different dimensions. All computations were carried out on a 2.3 GHz Quad-Core Intel i7 processor. The results show that even for 1000 dimensions, the problem is tractable for a Galerkin expansion in up to 6 dimensions on a laptop’s processor.
$n=2$ $n=100$ $n=1000$
time (s) 7.1 447 6935
Table 6: Table showing the computational time, in seconds, required to compute the approximate solution $\pi^{N}$ for $n=2,100,1000$, with $\Omega=[-1,1]^{2}$ and $M=6$.
Based on the computed $\pi^{N}$ for these different cases, it remains to use this approximate invariant mapping to construct a stable reduced-order model to mirror the output of the nonlinear full-order system. We recall from (13) that we would like a reduced-order model of the form
@@
where $\bar{g}$ is a free mapping. As long as $g(\cdot)$ is chosen so that the system is locally exponentially stable, the reduced-order model will achieve moment matching. Section 2.2 gave a way to construct this mapping in general, but for this particular case it can be shown that the mapping
@@
achieves asymptotic stability for some sufficiently large, positive constant $c$, and can therefore be used to construct the reduced-order model. This reduced-order model is given by

### v2

@@ 3.3 Integration
We are concerned with the assembly of $\boldsymbol{A}$, $G_{i}$, and $\frac{\partial G_{i}}{\partial\boldsymbol{c}_{j}}$. Each component of $\boldsymbol{A}=\left\langle\frac{\partial\Phi^{N}}{\partial\omega}s,\Phi^{N}\right\rangle_{\Omega}$ is given by
(42) $\boldsymbol{A}_{i,j}=\left\langle\frac{\partial\phi_{j}}{\partial\omega}s,\phi_{i}\right\rangle_{\Omega}=\sum_{k=1}^{d}\int_{\Omega}\frac{\partial\phi_{j}}{\partial\omega_{k}}(\omega)s_{k}(\omega)\phi_{i}(\omega)d\omega,$
where $s(\omega)=\left(s_{1}(\omega),\dots,s_{d}(\omega)\right)^{\top}$. When each $s_{k}(\omega)$ is a monomial, this expression can be computed exactly. Each component of $G_{i}(\boldsymbol{c})=-\left\langle f_{i}(\pi^{N},\ell),\Phi^{N}\right\rangle_{\Omega}$ is represented as
(43) $G_{i}(\boldsymbol{c})_{j}=-\left\langle f_{i}(\pi^{N},\ell),\phi_{j}\right\rangle_{\Omega}=-\int_{\Omega}f_{i}(\pi^{N}(\omega),\ell(\omega))\phi_{j}(\omega)d\omega,$
and is approximated using a suitable quadrature rule, for example Gauss-Legendre quadrature in $d$ dimensions over $\Omega$. The entries of $\frac{\partial G_{i}}{\partial\boldsymbol{c}_{j}}=-\left\langle\frac{\partial f_{i}(\pi^{N},\ell)}{\partial\boldsymbol{c}_{j}}\Phi^{N},\Phi^{N}\right\rangle_{\Omega}$ are computed analogously,
(44) ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\left[\frac{\partial G_{i}}{\partial\boldsymbol{c}_{j}}\right]}_{k,l}$ $=-\left\langle\frac{\partial f_{i}(\pi^{N},\ell)}{\partial\boldsymbol{c}_{j}}\phi_{l},\phi_{k}\right\rangle_{\Omega}=-\int_{\Omega}\frac{\partial f_{i}(\pi^{N}(\omega),\ell(\omega))}{\partial\boldsymbol{c}_{j}}\phi_{l}(\omega)\phi_{k}(\omega)d\omega,$
and are likewise approximated by a suitable quadrature rule.
@@ 3.4 Computational complexity and implementation
The bulk of the computation comes from integration over $d$ dimensions to evaluate the function $F$ and its Jacobian $JF$ as well as the computation of the Newton update $\Delta\boldsymbol{c}^{k}$. The computational complexity of integration scales exponentially with $d$ unless $f$ and $s$ are separable or have exploitable structure, as discussed in Remark 3.2; in such cases the integral can be handled via approximations such as the Monte Carlo method.
When the integrals admit an analytical solution, evaluating $F$ has complexity $O(Nn)$ and evaluating $JF$ has complexity $O(N^{2}n^{2})$. Solving $JF(\boldsymbol{c}^{k})\Delta\boldsymbol{c}^{k}=F(\boldsymbol{c}^{k})$ via a direct or least-squares solver has cubic complexity in $Nn$, though in many cases $JF$ is sparse, which significantly reduces this cost. Overall, the computational cost depends on the system dimension $n$ and combinatorially on the signal generator dimension $d$ and maximum degree $M$.
4 Numerical results
Different numerical experiments are performed to assess the approximation quality of the Galerkin method and the performance of the resulting reduced-order models. The tests are carried out over rectangular domains $\Omega$ of varying sizes, for different nonlinear dynamics and signal generators, and the algorithm is employed with a tolerance of $10^{-7}$ in the $L^{1}$ norm of $F$. In each case, the solver is initialized with $\boldsymbol{c}^{0}=\boldsymbol{0}$, with all coefficients of $\pi^{N}$ set to zero.
The quality of the Galerkin approximation is measured via the residual error function $R=(R_{1},\dots,R_{n})^{\top}$, where for each $i=1,\dots,n$ the residual at a specified $\omega$ is given by
(45) $R_{i}(\boldsymbol{c},\omega)=\frac{\partial\pi_{i}^{N}}{\partial\omega}(\omega)s(\omega)-f_{i}(\pi^{N}(\omega),\ell(\omega)).$
To quantify the approximation quality globally, we compute the $L_{2}$ norm of each $R_{i}$ averaged over a chosen test domain $W$,
(46) $\lVert R_{i}\rVert_{(2,W)}=\sqrt{\int_{W}|R_{i}(\boldsymbol{c},\omega)|^{2}d\omega},$
and aggregate these into a single metric weighted by the relative magnitude of the coefficients $\boldsymbol{c}_{i}$:
@@ 4 Numerical results
When a reduced-order model can be constructed, its performance is assessed by comparing steady-state outputs. Given time series $y_{k}=y(t_{k})$ and $y_{r,k}=y_{r}(t_{k})$ sampled at times $\{t_{k}\}_{k=1}^{K}$, we identify a time $T_{ss}$ beyond which both models have reached steady state, and set $k_{s}=\min\{k:t_{k}\geq T_{ss}\}$. Denoting by $\frac{1}{2}\big(\max_{k\geq k_{s}}y_{k}-\min_{k\geq k_{s}}y_{k}\big)$ the signal amplitude in steady state, the relative RMS error is
(48) $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\epsilon_{rel}=\frac{\big(\frac{1}{K_{s}}\sum_{k=k_{s}}^{K}\|y_{r,k}-y_{k}\|_{2}^{2}\big)^{1/2}}{\frac{1}{2}\big(\max_{k\geq k_{s}}y_{k}-\min_{k\geq k_{s}}y_{k}\big)},$}
where $K_{s}=K+1-k_{s}$ is the number of samples in the steady-state regime.
Three examples are considered. The first, in Section 4.1, serves as a proof of concept on a low-dimensional nonlinear system for which an analytical solution is known, demonstrating the algorithm’s ability to recover exact solutions when the components of $\pi$ belong to the span of the basis. The second and third examples, in Section 4.2 and Section 4.3, consider a high-dimensional system of dimension $n=1000$ driven by a linear and a nonlinear signal generator respectively, demonstrating scalability and the ability to construct accurate reduced-order models in the absence of an analytical solution.
4.1 Test 1: Cart pendulum
First, we consider the position control of a nonlinear cart pendulum, based on a problem studied in [32, Section 3.2]. The signal generator is chosen by selecting target dynamics for the cart pendulum with $d=2$ and $n=4$. The dynamics of the interconnected system are governed by
@@ 4.1 Test 1: Cart pendulum
This problem admits the analytical solution
@@
The function $F$ and its Jacobian $JF$ take the form
$F(\boldsymbol{c})=\begin{bmatrix}\boldsymbol{A}\boldsymbol{c}_{1}-\boldsymbol{M}\boldsymbol{c}_{3}\\ \boldsymbol{A}\boldsymbol{c}_{2}-\boldsymbol{M}\boldsymbol{c}_{4}\\ \boldsymbol{A}\boldsymbol{c}_{3}-H(\boldsymbol{c}_{1})\\ \boldsymbol{A}\boldsymbol{c}_{4}-\boldsymbol{\gamma}\end{bmatrix},\quad JF(\boldsymbol{c})=\begin{bmatrix}\boldsymbol{A}&0&-\boldsymbol{M}&0\\ 0&\boldsymbol{A}&0&-\boldsymbol{M}\\ -\frac{\partial H}{\partial\boldsymbol{c}_{1}}&0&\boldsymbol{A}&0\\ 0&0&0&\boldsymbol{A}\end{bmatrix}$
using the previously defined matrix $\boldsymbol{A}$ in (42), as well as the matrices and vectors
(49) $\boldsymbol{M}=\left\langle\Phi^{N},\Phi^{N}\right\rangle_{\Omega},$ $\boldsymbol{P}=\left\langle\Phi^{N}\omega_{1},\Phi^{N}\right\rangle_{\Omega},$ $\boldsymbol{\gamma}=\left\langle\ell(\omega),\Phi^{N}\right\rangle_{\Omega},$
and the function $H:\mathbb{R}^{N}\to\mathbb{R}^{N}$ given by
@@
Due to the choice of $s(\omega)$, the matrix $\boldsymbol{A}$ is singular and consequently $JF$ is also singular; Newton iteration updates are therefore computed in the least-squares sense. Using the constants $a_{1}=2$, $a_{2}=3$, and $k=-\frac{2}{3}$, Table 1 shows that the method achieves near-exact recovery of the analytical solution over $\Omega=[-1,1]^{2}$. The solution admits a monomial representation and the algorithm recovers it to within the prescribed numerical tolerance, with observed errors on the order of $10^{-9}$.
$M=2$ $M=4$ $M=6$ $M=8$
$\|R(\mathbf{c},\omega)\|_{(2,W)}$ $9.3\times 10^{-10}$ $8.6\times 10^{-10}$ $9.4\times 10^{-10}$ $2.4\times 10^{-9}$
Time (s) $0.83$ $16.8$ $135$ $646$
Table 1: Residual error $\lVert R(\boldsymbol{c},\omega)\rVert_{(2,W)}$ and computational time for the Galerkin approximation of the invariant mapping $\pi$ for Test 1.
4.2 Test 2: Nonlinear resistor-inductor ladder with linear oscillator as signal generator
We consider the nonlinear resistor-inductor (RL) ladder, based on a slightly modified version of the circuit schematic in [34, Figure 7] and also considered in [24] in a model order reduction context. The signal generator has dimension $d=2$ and the system has a repeating structure admitting any state dimension $n$. The signal generator is given by
@@ 4.2 Test 2: Nonlinear resistor-inductor ladder with linear oscillator as signal generator
The parameter $\kappa=1$ renders the system unstable beyond a certain dimension $n$; we therefore use $\kappa=1.1$ throughout. Since the nonlinearities in each $f_{i}$ are polynomials of degree 2 and 3 depending only on $x_{i}$, the construction of $F$ and $JF$ admits the following simplifications. We introduce the tensors
(52) $\boldsymbol{N}\in\mathbb{R}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}N\times N\times N}},[\boldsymbol{N}]_{i,j,k}=\int_{\Omega}\phi_{i}\phi_{j}\phi_{k}d\omega$
$\boldsymbol{O}\in\mathbb{R}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}N\times N\times N\times N}},[\boldsymbol{O}]_{i,j,k,l}=\int_{\Omega}\phi_{i}\phi_{j}\phi_{k}\phi_{l}d\omega,$
and the matrix-valued functions ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{N}}:\mathbb{R}^{N}\to\mathbb{R}^{N\times N}$ and ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{O}}:\mathbb{R}^{N}\to\mathbb{R}^{N\times N}$,
(53) ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\left[\mathcal{N}(\boldsymbol{v})\right]_{i,j}}=\sum_{k=1}^{N}v_{k}N_{i,j,k},$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\left[\mathcal{O}(\boldsymbol{v})\right]_{i,j}}=\sum_{k=1}^{N}\sum_{l=1}^{N}v_{k}v_{l}O_{i,j,k,l}.$
Defining $P,Q:\mathbb{R}^{N}\to\mathbb{R}^{N\times N}$ as
(54) $P(\boldsymbol{v})$ $=\left(\boldsymbol{A}+2\kappa\boldsymbol{M}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{1}{2}\mathcal{N}(\boldsymbol{v})+\frac{1}{3}\mathcal{O}(\boldsymbol{v})}\right)\boldsymbol{v}$
$Q(\boldsymbol{v})$ $=\boldsymbol{A}+2\kappa\boldsymbol{M}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{N}(\boldsymbol{v})+\mathcal{O}(\boldsymbol{v})},$
where $\boldsymbol{A}$ and $\boldsymbol{M}$ are as in (42) and (49), the function $F$ and its Jacobian take the form
@@
where $\boldsymbol{\gamma}$ is the same as in (49), and $\delta_{i,j}$ denotes the Kronecker delta, and
@@
Computing the tensors $\boldsymbol{N}$ and $\boldsymbol{O}$ exactly rather than by numerical quadrature allows for a significant reduction in computational cost by exploiting the structure of the system dynamics.
The method is validated on a full-order model with $n=1000$ and $a=2$, over the domain $\Omega=[-0.6,0.6]^{2}$, with residual errors measured on $W=\Omega$. The choice of domain is discussed further in Section 5.1. The residual errors $\lVert R(\boldsymbol{c},\omega)\rVert_{(2,W)}$ reported in Table 2 show that increasing $M$ systematically improves the approximation, as expected from the nested structure of the polynomial spaces. The small residual errors at higher degrees suggest that $0.6,0.6]^{2}$ lies within the neighbourhood $W^{0}$ over which $\pi$ is well-defined, and that $\pi^{N}$ converges to the weak solution of the invariance equation as the number of basis functions increases.
To construct the reduced-order model, we recall from (17) the family of models
@@
[…]
@@
achieves local exponential stability for any $c>0$. To verify this, we examine the linearization about the origin of $\frac{\partial}{\partial r}\left(s(r)-\begin{bmatrix}0\\ c\end{bmatrix}\ell(r)\right)$, whose eigenvalues are those of $\begin{bmatrix}0&a\\ -a&-c\end{bmatrix}$, which have strictly negative real part for all $c>0$. The reduced-order model is then